Hi there,
I can't believe I just don't get this problem. It's killing me and I can't go to sleep without solving it. Any help would be appreciated!
The complement of an angle measures 10 Degrees less than one-fifth of its supplement. Find the measure of the angle.
THANK YOU!!!!

Found 2 solutions by edjones, Earlsdon:
Answer by edjones(8007)

Let a=the angle
The complement=90-a
the supplement is 180-a
(90-a)+10=(180-a)/5
5(100-a)=180-a
500-5a=180-a
320=4a
a=80 degrees
.
Ed

Answer by Earlsdon(6294)

Let the angle be A, then:
90-A = (1/5)(180-A)-10 Simplify and solve for A.
90-A = 36-(1/5)A-10
90-A = 26-(1/5)A Add A to both sides.
90 = 26+(4/5)A Subtract 26 from both sides.
64 = (4/5)A Multiply both sides by 5/4
80 = A
The angle is 80 degrees.
Check:
90-80 = (1/5)(180-80)-10
10 = (1/5)(100)-10
10 = 20-10
10 = 10
Good night and sleep well!